# Katsura’s self-similar groupoid actions, Putnam’s binary factors, and their limit spaces

### v1

@@ 4.2. Putnam’s binary factor maps as self-similar groupoid actions on graphs
$(G_{\xi},E)$ is not a self-similar groupoid action arising from a Katsura groupoid. Indeed, if $(G_{\xi},E)$ was isomorphic to $(G_{B},E_{A})$ for some Katsura pair, then $A=(3)$ and $B=(n)$ for some $n\in\mathbb{Z}$. The action of $G_{\xi}=\mathbb{Z}$ on $E$ is non-trivial and contracting, so its contracting coefficient satisfies $10<\rho=\frac{|n|}{3}<$ 1. So either $|n|=1$ or $|n|=2$, but neither of these cases allow for $\mathbb{Z}\cdot f=f$, $\mathbb{Z}|_{f}=0$. Thus, $(G_{\xi},E)$ is not isomorphic to a Katsura groupoid.
However, we will show in Section 5 that every self-similar groupoid from an embedding pair is an out-splitting of a Katsura groupoid. ∎
@@ 4.3. Properties of $(G_{\xi},E)$
5. Out-splittings and Katsura groupoid models for embedding pairs
In this section, we determine the relationship of $(G_{\xi},E)$ with a Katsura groupoid model. In particular, we show there are matrices $A\in M_{N}(\{0,1,2\})$ and $B\in M_{N}(\{0,1\})$ such that $(\tilde{\sigma},\mathcal{J}_{G_{\xi},E})$ is topologically conjugate to $(\tilde{\sigma},\mathcal{J}_{G_{B},E_{A}})$. We will do so by showing out-splits of self-similar group bundles preserve limit spaces, and that a certain Katsura groupoid arises as an out-split of $(G_{\xi},E)$.
@@ 5.1. Out-splits of self-similar group bundles
The action-restriction pair $(G_{OS},E_{OS})$ is defined for $(g,v)\in G_{OS}$ and $(v,e)\in G_{OS}^{1}$, as
@@
5.2. Katsura groupoid models for embedding pairs
@@
Let $E_{OS}^{0}$ be the set obtained from $E^{1}$ by identifying the edges $\xi^{0}(h)$ and $\xi^{1}(h)$, for all $h\in H^{1}$. Let $\pi:E^{1}\to E_{OS}^{0}$ be the quotient map. Since the edges being identified share the same source, there is a unique map $\beta:E_{OS}^{0}\to E^{0}$ satisfying $\beta\circ\pi=s$. Therefore, the tuple $OS=(\pi,\beta)$ determines an out-split $(G_{OS},E_{OS})$ of the self-similar group bundle $(G_{\xi},E)$. We show $(G_{OS},E_{OS})$ is isomorphic to a Katsura groupoid action.
@@
Let $\xi=\xi^{0},\xi^{1}:H\to E$ be an embedding pair and let $(G_{\xi},E)$ be its associated self-similar groupoid action, as described in Section 4.2. Then, the out-split $(G_{OS},E_{OS})$ of $(G_{\xi},E)$, described in Section 5.2 is canonically isomorphic to the Katsura groupoid $(G_{B},E_{A})$ with $A=(A_{v,w})_{v,w\in E_{OS}^{0}}$ the adjacency matrix of $E_{OS}$ and $B=(\max\{0,A_{v,w}-1\})_{v,w\in E_{OS}^{0}}$. Moreover, the limit spaces $(\tilde{\sigma},\mathcal{J}_{G_{\xi},E})$ and $(\tilde{\sigma},\mathcal{J}_{G_{B},E_{A}})$ are topologically conjugate.
@@
$E_{OS}^{1}=\{(f,f),(e,f),(e,e_{i}),(f,e_{j})\mid i,j\in\{1,2\}\}.$
The graph $E$ and its out-split $OS$ is recorded in Figure 4. If we order $e<f$, then the Katsura groupoid is defined by the matrices
@@
6. Katsura systems as bundles of odometers
We now provide a description of the Katsura systems $(\tilde{\sigma},\mathcal{J}_{G_{B},E_{A}})$ as a bundle of dynamical systems that fibre over the shift space of the connectivity graph of $E_{A}$ when $B$ is a matrix taking values in $\{0,1\}$. Note that from the previous section, this class includes the dynamical systems arising from embedding pairs.
@@ 6. Katsura systems as bundles of odometers
Let $(G_{B},E_{A})$ be a Katsura groupoid such that $B\in M_{N}(\{0,1\})$. For $\mu,\nu\in E^{-\infty}_{A}$, $\mu\sim_{e}\nu$ if and only if $\mu=\nu$, or there is $K<0$ such that $s(\mu_{k})=s(\nu_{k})\coloneqq v_{k}$ and $B_{v_{k-1},v_{k}}=1$ for all $k\leq K$, and $B_{v_{K},v_{K+1}}=0$, $\mu_{K+1}...\mu_{-1}=\nu_{K+1}...\nu_{-1}$ if $K<1$.
@@
Since the action of a Katsura groupoid preserves the vertices of paths, there are factor maps $\pi_{\mathcal{J}}:\mathcal{J}_{G_{B},E_{A}}\to E^{-\infty}_{C}$ and $\pi_{\mathcal{C}}:\mathcal{C}_{G_{B},E_{A}}\to E^{-\infty}_{C}$, where $C$ is the connectivity matrix of $E_{A}$. We can use this factor map to describe the connected components of $\mathcal{J}_{G_{B},E_{A}}$. The following fact is contained in the proof of Proposition 6.3.
@@
Let $(G_{B},E_{A})$ be a Katsura groupoid such that $B\in M_{N}(\{0,1\})$ and $z\in\mathcal{J}_{(G_{B},E_{A})}$. Denote $\pi_{\mathcal{J}}(z)=v$. Then, $z$ is a connected component if $B_{v_{k-1},v_{k}}=0$ infinitely often.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid such that $B\in M_{N}(\{0,1\})$. Suppose $v\in E_{C}^{-\infty}$ satisfies $B_{v_{k-1},v_{k}}=1$ for all $k\in\mathbb{N}$. For $\mu,\nu\in\mathcal{A}_{v}$, $\mu\sim_{ae}\nu$ if and only if $C_{v}(\mu)=C_{v}(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid such that $B\in M_{N}(\{0,1\})$. Then, a connected component of $\mathcal{J}_{G_{B},E_{A}}$ is either a point or a circle.
@@ 7. Planar embedding of Putnam’s spaces
In [13, Question 7.10], Putnam asked when $\mathcal{J}_{\xi}=E^{-\infty}/\sim_{\xi}$ embeds into the plane. In this section, we prove $\mathcal{J}_{\xi}$ always embeds into the plane by using our description of $\mathcal{J}_{\xi}$ as the limit space of a regular Katsura groupoid. In particular, Theorem 6.6 gave a topological description of the limit space when $B\in M_{N}(\{0,1\})$ as a Cantor set bundle of circles (with the convention that a point is a circle of radius zero). We use this as inspiration to define an embedding from the limit space to the complex plane whenever $B\in M_{N}(\{0,1\})$ and $(G_{B},E_{A})$ is regular. See Corollary 3.17 for a characterisation of regularity in terms of $A$ and $B$. Notice that Section 5.2 proved that the Katsura groupoids from embedding pairs always have the property that $A\in M_{N}(\{0,1,2\})$ and $B=(B_{ij})$ where $B_{ij}=\max\{A_{ij}-1,0\}$. Thus, Putnam’s question will be answered by the following more general result.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid action such that $B\in M_{N}(\{0,1\})$. Then, there is a continuous injection $\zeta:\mathcal{J}_{G_{B},E_{A}}\to\mathbb{C}$.
@@
Let $(G_{B},E_{A})$ be the Katsura groupoid action defined by
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid action such that $B\in M_{N}(\{0,1\})$. If $\mu,\nu\in E_{A}^{-\infty}$ satisfy $\mu\sim_{ae}\nu$, then $\zeta(\mu)=\zeta(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid action such that $B\in M_{N}(\{0,1\})$. Suppose $\mu,\nu$ are in $E^{-\infty}_{A}$ and let $J\in\mathcal{I}(\mu)$ and $K\in\mathcal{I}(\nu)$ be the intervals containing $-1$, respectively. If $J^{-}\leq K^{-}$, $-\infty<K^{-}$ and $\Omega(\mu_{K})\neq\Omega(\nu_{K})$, then $\zeta(\mu)\neq\zeta(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid action such that $B\in M_{N}(\{0,1\})$. Suppose $\mu,\nu$ are in $E^{-\infty}_{A}$ and let $J\in\mathcal{I}(\mu)$ and $K\in\mathcal{I}(\nu)$ be the intervals containing $-1$, respectively. If $J^{-}<K^{-}$, then $\zeta(\mu)\neq\zeta(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid action such that $B\in M_{N}(\{0,1\})$. Suppose $\mu\neq\nu$ are in $E^{-\infty}_{A}$ such that $\zeta(\mu)=\zeta(\nu)$. Let $j<0$ be the largest number such that $\mu_{j}\neq\nu_{j}$ and let $J_{1}(\mu)\in\mathcal{I}(\mu)$ and $J_{1}(\nu)\in\mathcal{I}(\nu)$ be the intervals containing $j$. Then $J_{1}(\mu)^{+}=J_{1}(\nu)^{+}$.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid action such that $B$ is in $M_{N}(\{0,1\})$. Suppose $\mu,\nu$ are in $E^{-\infty}_{A}$ and let $J\in\mathcal{I}(\mu)$ and $K\in\mathcal{I}(\nu)$ be the intervals containing $-1$, respectively. If $-\infty<J^{-}$, $J=K$, and $\Omega(\mu_{J})=\Omega(\nu_{J})$, then $\theta^{k}(\mu_{J})\neq\theta^{k}(\nu_{J})$ implies $\zeta(\mu)\neq\zeta(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid action such that $B\in M_{N}(\{0,1\})$. If $\mu,\nu\in E^{-\infty}_{A}$ are such that $\zeta(\mu)=\zeta(\nu)$, then either $\mu=\nu$ or there is $k<0$ such that $\mu_{[k,-1]}=\nu_{[k,-1]}$, $[-\infty,k-1]\in\mathcal{I}^{1}(\mu)\cap\mathcal{I}^{1}(\nu)$, $s(\mu_{j})=s(\nu_{j})$ for all $j\leq k-1$ and $\theta^{1}(\mu_{[-\infty,k-1]})=\theta^{1}(\nu_{[-\infty,k-1]})$.
@@
Let $(G_{B},E_{A})$ be a regular Katsura groupoid action such that $B\in M_{N}(\{0,1\})$. If $\mu,\nu\in E^{-\infty}_{A}$ are such that $\zeta(\mu)=\zeta(\nu)$, then $\mu\sim_{ae}\nu$. Therefore, $\zeta:\mathcal{J}_{G_{B},E_{A}}\to\mathbb{C}$ is an embedding.

### v2

@@ 4.2. Putnam’s binary factor maps as self-similar groupoid actions on graphs
$(G_{\xi},E)$ is not a self-similar groupoid action arising from a KEP-action. Indeed, if $(G_{\xi},E)$ was isomorphic to $(G_{B},E_{A})$ for some Katsura pair, then $A=(3)$ and $B=(n)$ for some $n\in\mathbb{Z}$. The action of $G_{\xi}=\mathbb{Z}$ on $E$ is non-trivial and contracting, so its contracting coefficient satisfies $10<\rho=\frac{|n|}{3}<$ 1. So either $|n|=1$ or $|n|=2$, but neither of these cases allow for $\mathbb{Z}\cdot f=f$, $\mathbb{Z}|_{f}=0$. Thus, $(G_{\xi},E)$ is not isomorphic to a KEP-action.
However, we will show in Section 5 that every self-similar groupoid from an embedding pair is an out-splitting of a KEP-action. ∎
@@ 4.3. Properties of $(G_{\xi},E)$
5. Out-splittings and KEP-action models for embedding pairs
In this section, we determine the relationship of $(G_{\xi},E)$ with a KEP-action model. In particular, we show there are matrices $A\in M_{N}(\{0,1,2\})$ and $B\in M_{N}(\{0,1\})$ such that $(\tilde{\sigma},\mathcal{J}_{G_{\xi},E})$ is topologically conjugate to $(\tilde{\sigma},\mathcal{J}_{G_{B},E_{A}})$. We will do so by showing out-splits of self-similar group bundles preserve limit spaces, and that a certain KEP-action arises as an out-split of $(G_{\xi},E)$.
@@ 5.1. Out-splits of self-similar group bundles
The action-restriction pair $(G_{OS},E_{OS})$ is defined for $(g,v)\in G_{OS}$ and $(v,e)\in E_{OS}^{1}$, as
@@
5.2. KEP-action models for embedding pairs
@@
Let $E_{OS}^{0}$ be the set obtained from $E^{1}$ by identifying the edges $\xi^{0}(h)$ and $\xi^{1}(h)$, for all $h\in H^{1}$. Let $\pi:E^{1}\to E_{OS}^{0}$ be the quotient map. Since the edges being identified share the same source, there is a unique map $\beta:E_{OS}^{0}\to E^{0}$ satisfying $\beta\circ\pi=s$. Therefore, the tuple $OS=(\pi,\beta)$ determines an out-split $(G_{OS},E_{OS})$ of the self-similar group bundle $(G_{\xi},E)$. We show $(G_{OS},E_{OS})$ is isomorphic to a KEP-action.
@@
Let $\xi=\xi^{0},\xi^{1}:H\to E$ be an embedding pair and let $(G_{\xi},E)$ be its associated self-similar groupoid action, as described in Section 4.2. Then, the out-split $(G_{OS},E_{OS})$ of $(G_{\xi},E)$, described in Section 5.2 is canonically isomorphic to the KEP-action $(G_{B},E_{A})$ with $A=(A_{v,w})_{v,w\in E_{OS}^{0}}$ the adjacency matrix of $E_{OS}$ and $B=(\max\{0,A_{v,w}-1\})_{v,w\in E_{OS}^{0}}$. Moreover, the limit spaces $(\tilde{\sigma},\mathcal{J}_{G_{\xi},E})$ and $(\tilde{\sigma},\mathcal{J}_{G_{B},E_{A}})$ are topologically conjugate.
@@
$E_{OS}^{1}=\{(f,f),(e,f),(e,e_{i}),(f,e_{j})\mid i,j\in\{0,1\}\}.$
The graph $E$ and its out-split $OS$ is recorded in Figure 4. If we order $e<f$, then the KEP-action is defined by the matrices
@@
6. KEP-systems as bundles of odometers
We now provide a description of the KEP-systems $(\tilde{\sigma},\mathcal{J}_{G_{B},E_{A}})$ as a bundle of dynamical systems that fibre over the shift space of the connectivity graph of $E_{A}$ when $B$ is a matrix taking values in $\{0,1\}$. Note that from the previous section, this class includes the dynamical systems arising from embedding pairs.
@@ 6. KEP-systems as bundles of odometers
Let $(G_{B},E_{A})$ be a KEP-action such that $B\in M_{N}(\{0,1\})$. For $\mu,\nu\in E^{-\infty}_{A}$, $\mu\sim_{e}\nu$ if and only if $\mu=\nu$, or there is $K<0$ such that $s(\mu_{k})=s(\nu_{k})\coloneqq v_{k}$ and $B_{v_{k-1},v_{k}}=1$ for all $k\leq K$, and $B_{v_{K},v_{K+1}}=0$, $\mu_{K+1}...\mu_{-1}=\nu_{K+1}...\nu_{-1}$ if $K<1$.
@@
Since the action of a KEP-action preserves the vertices of paths, there are factor maps $\pi_{\mathcal{J}}:\mathcal{J}_{G_{B},E_{A}}\to E^{-\infty}_{C}$ and $\pi_{\mathcal{C}}:\mathcal{C}_{G_{B},E_{A}}\to E^{-\infty}_{C}$, where $C$ is the connectivity matrix of $E_{A}$. We can use this factor map to describe the connected components of $\mathcal{J}_{G_{B},E_{A}}$. The following fact is contained in the proof of Proposition 6.3.
@@
Let $(G_{B},E_{A})$ be a KEP-action such that $B\in M_{N}(\{0,1\})$ and $z\in\mathcal{J}_{(G_{B},E_{A})}$. Denote $\pi_{\mathcal{J}}(z)=v$. Then, $z$ is a connected component if $B_{v_{k-1},v_{k}}=0$ infinitely often.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. Suppose $v\in E_{C}^{-\infty}$ satisfies $B_{v_{k-1},v_{k}}=1$ for all $k\in\mathbb{N}$. For $\mu,\nu\in\mathcal{A}_{v}$, $\mu\sim_{ae}\nu$ if and only if $C_{v}(\mu)=C_{v}(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. Then, a connected component of $\mathcal{J}_{G_{B},E_{A}}$ is either a point or a circle.
@@ 7. Planar embedding of Putnam’s spaces
In [13, Question 7.10], Putnam asked when $\mathcal{J}_{\xi}=E^{-\infty}/\sim_{\xi}$ embeds into the plane. In this section, we prove $\mathcal{J}_{\xi}$ always embeds into the plane by using our description of $\mathcal{J}_{\xi}$ as the limit space of a regular KEP-action. In particular, Theorem 6.6 gave a topological description of the limit space when $B\in M_{N}(\{0,1\})$ as a Cantor set bundle of circles (with the convention that a point is a circle of radius zero). We use this as inspiration to define an embedding from the limit space to the complex plane whenever $B\in M_{N}(\{0,1\})$ and $(G_{B},E_{A})$ is regular. See Corollary 3.14 for a characterisation of regularity in terms of $A$ and $B$. Notice that Section 5.2 proved that the KEP-actions from embedding pairs always have the property that $A\in M_{N}(\{0,1,2\})$ and $B=(B_{ij})$ where $B_{ij}=\max\{A_{ij}-1,0\}$. Thus, Putnam’s question will be answered by the following more general result.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. Then, there is a continuous injection $\zeta:\mathcal{J}_{G_{B},E_{A}}\to\mathbb{C}$.
@@
Let $(G_{B},E_{A})$ be the KEP-action defined by
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. If $\mu,\nu\in E_{A}^{-\infty}$ satisfy $\mu\sim_{ae}\nu$, then $\zeta(\mu)=\zeta(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. Suppose $\mu,\nu$ are in $E^{-\infty}_{A}$ and let $J\in\mathcal{I}(\mu)$ and $K\in\mathcal{I}(\nu)$ be the intervals containing $-1$, respectively. If $J^{-}\leq K^{-}$, $-\infty<K^{-}$ and $\Omega(\mu_{K})\neq\Omega(\nu_{K})$, then $\zeta(\mu)\neq\zeta(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. Suppose $\mu,\nu$ are in $E^{-\infty}_{A}$ and let $J\in\mathcal{I}(\mu)$ and $K\in\mathcal{I}(\nu)$ be the intervals containing $-1$, respectively. If $J^{-}<K^{-}$, then $\zeta(\mu)\neq\zeta(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. Suppose $\mu\neq\nu$ are in $E^{-\infty}_{A}$ such that $\zeta(\mu)=\zeta(\nu)$. Let $j<0$ be the largest number such that $\mu_{j}\neq\nu_{j}$ and let $J_{1}(\mu)\in\mathcal{I}(\mu)$ and $J_{1}(\nu)\in\mathcal{I}(\nu)$ be the intervals containing $j$. Then $J_{1}(\mu)^{+}=J_{1}(\nu)^{+}$.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B$ is in $M_{N}(\{0,1\})$. Suppose $\mu,\nu$ are in $E^{-\infty}_{A}$ and let $J\in\mathcal{I}(\mu)$ and $K\in\mathcal{I}(\nu)$ be the intervals containing $-1$, respectively. If $-\infty<J^{-}$, $J=K$, and $\Omega(\mu_{J})=\Omega(\nu_{J})$, then $\theta^{k}(\mu_{J})\neq\theta^{k}(\nu_{J})$ implies $\zeta(\mu)\neq\zeta(\nu)$.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. If $\mu,\nu\in E^{-\infty}_{A}$ are such that $\zeta(\mu)=\zeta(\nu)$, then either $\mu=\nu$ or there is $k<0$ such that $\mu_{[k,-1]}=\nu_{[k,-1]}$, $[-\infty,k-1]\in\mathcal{I}^{1}(\mu)\cap\mathcal{I}^{1}(\nu)$, $s(\mu_{j})=s(\nu_{j})$ for all $j\leq k-1$ and $\theta^{1}(\mu_{[-\infty,k-1]})=\theta^{1}(\nu_{[-\infty,k-1]})$.
@@
Let $(G_{B},E_{A})$ be a regular KEP-action such that $B\in M_{N}(\{0,1\})$. If $\mu,\nu\in E^{-\infty}_{A}$ are such that $\zeta(\mu)=\zeta(\nu)$, then $\mu\sim_{ae}\nu$. Therefore, $\zeta:\mathcal{J}_{G_{B},E_{A}}\to\mathbb{C}$ is an embedding.